10x+5 i need help with factoring
Answer by Jessmh123(22)

lets say one book is x and the other is y
we know that x+y=390
one book is 40 more than 4 times the other, we can write that like like...
x=4y+40
now that we know what x is equal to, we can plug it back into our first equation, which was x+y=390.
so...
4y+40+y=390
5y+40=390 (combime like terms)
5y=350 (subtract 40 from each side)
y=70 (divide by 5 on both sides to solve for y)
now we know that one book has 70 pages, and now we can just subtract 70 from 390 to find out how many pages the other book has
390-70=320
one books had 70 pages and the other has 320
:)